Together a 100 cm wide escalator and a 60 cm wide escalator can empty a 1575 person room in 14 minutes. The wider escalator moves twice as many people as the narrow escalator. How many people per hour does the 60 cm wide escalator move?
Answer by rfer(16322)

small escalator=x
large escalator=2x
14 min=14/60
x+2x=1575/(14/60)
3x=6750
x=2250 small escalator
2x=4500 large escalator